# Forecasting causal dynamics with universal reservoirs

### v1

@@ 4 Numerical illustration
Figure 1: Data from a trajectory used for training and forecasting the Lorenz system.
Figure 2: Log prediction error for the Lorenz system with full access to system states, averaged over 10 trajectories.
To learn the Lorenz system, we chose an ESN with $1000$ neurons whose connectivity matrix had spectral radius $1.2$, and whose input mask was scaled by $\gamma=13/3$. Inputs were taken as the states of the entire system, sampled at a time step of $\Delta t=0.02$ over $10000$ time steps. A washout of $1000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.5}$, between the state space trajectory and desired output trajectory on the remaining $9000$ time steps, the readout $h$ was trained and predictions made for an additional $2500$ time steps. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=0.9796$ using Bennetin’s algorithm over the length of one trajectory, and the multiplicative constant in the error bound (3.14) was estimated by fitting the predicted error bound to the actual errors. Plotted in Figure 2 are the log errors in prediction over $13$ Lyapunov times for an RC trained over $10$ trajectories with initial conditions chosen from a Gaussian distribution around $(0,1,1.05)$, averaged at each point across the $10$ trajectories. The log errors grow approximately at a linear rate until saturation of the attractor around $10$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
Figure 3: Predictions of a single trajectory for an ESN trained on the first coordinate of the Lorenz system.
Figure 4: Log prediction error for Lorenz system with access to the first coordinate, averaged over 10 trajectories.
@@
To illustrate our result for a causal chain, we took the observation function $\omega\colon\mathbb{R}^{3}\longrightarrow\mathbb{R},(x,y,z)\mapsto x,$ on the Lorenz system. By Corollary 2.2, sequences of observations through $\omega$ form a causal chain. To learn the Lorenz system using inputs only from the first coordinate, we chose an ESN with $1000$ neurons whose connectivity matrix had spectral radius $1.0$, and whose input mask was scaled by $\gamma=2.4$. Inputs were sampled at a time step of $\Delta t=0.005$ over $12000$ time steps. A washout of $1000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.75}$, between the state space trajectory and desired output trajectory on the remaining $11000$ time steps, the readout $h$ was trained and predictions made for an additional $4000$ time steps. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=0.9978$ using Bennetin’s algorithm over the length of one trajectory, and the multiplicative constant in the error bound (3.14) was estimated by fitting the predicted error bound to the actual errors. We chose $10$ initial conditions from a Gaussian distribution around $(0,1,1.05)$ and trained the readout of the RC over a single trajectory. In Figure 4 the log errors in prediction averaged across the $10$ trajectories are plotted for $10$ Lyapunov times. The errors grow approximately at a linear rate until saturation of the attractor around $7$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
@@
Figure 5: Log prediction error for the Rössler system with access to full system state, averaged over 10 trajectories.
To learn the Rössler system, we chose an ESN with $1000$ neurons and connectivity matrix with spectral radius $0.2$. The input mask was scaled by $\gamma=4.65$. Inputs were taken as the states of the entire system, sampled at a time step of $\Delta t=0.005$ over $23000$ time steps. A washout of $2000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.5}$, between the state space trajectory and desired output trajectory on the remaining $21000$ time steps, the readout $h$ was trained and predictions made for an additional $30000$ time steps. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=0.0814$ using Bennetin’s algorithm over the length of one trajectory, and the multiplicative constant in the error bound (3.14) was estimated by fitting the predicted error bound to the actual errors. We chose $10$ initial conditions from a Gaussian distribution around $(0.1,0.1,14)$ and trained the readout of the RC over a single trajectory. In Figure 5 the log errors in prediction averaged across the $10$ trajectories are plotted for $10$ Lyapunov times. The errors grow approximately at a linear rate until saturation of the attractor around $8$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
@@ Discussion.
The numerical experiments illustrate how the error committed using an ESN to predict three different systems grows at an exponential rate, governed by the top Lyapunov exponent associated with the learned ESN. The interplay between the Lyapunov exponents of the dynamical system and the ensuing effect on forecasting accuracy can clearly be seen in the numerical examples illustrated above. In all three examples the Lyapunov exponents of the learnt dynamical system associated to the trained RC are close to those of the underlying dynamical system being learnt. The Lyapunov exponent of the Lorenz system is greater than that of the Rössler system, leading to a quick saturation of the prediction error, in comparison with the RC trained on the Rössler system, which exhibits the ability to maintain reasonably accurate predictions long into the future. Nevertheless, scaling the time axis in Lyapunov times, shows that the prediction capabilities of ESNs on the two different systems are comparable.
It is to be noted, however, that in cases where the RC does not exhibit accurate long term predictions of the given system, the bound is not as visually apparent, or may indeed be redundant due to the swift saturation of the attractor. In this case errors settle at a finite value, determined by the diameter and stationary measure of the attractor, before the bound comes into play. This may be seen in part in the comparison between predicting the Lorenz system with access to full dynamical states and with access only to a single coordinate. The latter task is more challenging, due to restricted access to information about the system. Thus, while both RCs display similar Lyapunov exponents, the forecasting error in predicting the causal chain reaches saturation before that in predicting the entire system.
@@ 5 Conclusions
As noted in the numerical section, the bounds given have a number of limitations. In particular, it is not clear how to calculate an important multiplicative constant to fit the error bound to real case prediction scenarios. Nevertheless, in the case of well-trained state space systems used in prediction, the bound demonstrates a striking sharpness, accurately predicting the growth rate of the true forecasting error.

### v2

@@ 4 Numerical illustration
Figure 1: Data from a trajectory used for training and forecasting the Lorenz system with access to full coordinatees. The first black line divides the washout from the trajectory data used in training the ESN. The second black line divides the training trajectory from the trajectory used to test forecasting.
Figure 2: Prediction error for the Lorenz system with full access to system states, averaged over 100 trajectories. The error bound plotted has been averaged over the 100 trajectories. The gray band is an emperically calculated $95\%$ confidence interval for the errors committed by the $100$ trajectories. These are all plotted in the log scale.
To learn the Lorenz system, we chose an ESN with $1000$ neurons whose connectivity matrix had spectral radius $1.2$, and whose input mask was scaled by $\gamma=13/3$. Inputs were taken as the states of the entire system, sampled at a time step of $\Delta t=0.02$ over $10000$ time steps. A washout of $1000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.5}$, between the state space trajectory and desired output trajectory on the remaining $9000$ time steps, the readout $h$ was trained and predictions made for an additional $2500$ time steps. Plotted in Figure 2 are the errors in prediction over $10$ Lyapunov times for an RC trained over $100$ trajectories with initial conditions chosen from a Gaussian distribution around $(0,1,1.05)$ with standard deviation $1\times 10^{-10}$, averaged at each point across the $100$ trajectories and plotted in the log scale. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=0.8733$ using Bennetin’s algorithm over the length of one trajectory. The value of $\delta$ was fixed at $10^{-3}$ and $\varepsilon=8.988\times 10^{-6}$ was calculated by taking the maximum one-step ahead prediction error of the trained ESN over the training portions of the $100$ trajectories. The constant $R$ from (3.12) was calculated for each trajectory using (4.5) and then averaged. Figure 2 shows the error bound using this average value, plotted in the log scale. The log errors grow approximately at a linear rate until saturation of the attractor shortly before $9$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
Figure 3: Predictions and errors for a single trajectory of an ESN trained on the first coordinate of the Lorenz system.
Figure 4: Prediction error for the Lorenz system with access to the first coordinate, averaged over 100 trajectories. The error bound plotted has been averaged over the 100 trajectories. The gray band is an emperically calculated $95\%$ confidence interval for the errors committed by the $100$ trajectories. These are all plotted in the log scale.
@@
To illustrate our result for a causal chain, we took the observation function $\omega\colon\mathbb{R}^{3}\longrightarrow\mathbb{R},(x,y,z)\mapsto x$, on the Lorenz system. By Corollary 2.2, sequences of observations through $\omega$ form a causal chain. To learn the Lorenz system using inputs only from the first coordinate, we chose an ESN with $1000$ neurons whose connectivity matrix had spectral radius $1.0$, and whose input mask was scaled by $\gamma=2.4$. Inputs were sampled at a time step of $\Delta t=0.005$ over $12000$ time steps. A washout of $1000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.75}$, between the state space trajectory and desired output trajectory on the remaining $11000$ time steps, the readout $h$ was trained and predictions made for an additional $4000$ time steps. Plotted in Figure 4 are the errors in prediction over $9$ Lyapunov times for an RC trained on a single trajectory and giving predictions over $100$ trajectories with initial conditions chosen from a Gaussian distribution around $(0,1,1.05)$ with standard deviation $1\times 10^{-10}$, averaged at each point across the $100$ trajectories and plotted in the log scale. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=1.015$ using Bennetin’s algorithm over the length of one trajectory. The value of $\delta$ was fixed at $10^{-3}$ and $\varepsilon=5.733\times 10^{-7}$ was calculated by taking the maximum one-step ahead prediction error of the trained ESN over the training portions of the $100$ trajectories. The constant $R$ from (3.12) was calculated for each trajectory using (4.5) and then averaged. Figure 4 shows the error bound using this average value, plotted in the log scale. The log errors grow approximately at a linear rate until saturation of the attractor around $7$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
Figure 3 shows the forecasting errors and fitted bound for the $24^{th}$ trajectory from our dataset. The ESN exhibited good learning for this trajectory, giving accurate predictions for about $8$ Lyapunov times. As such it provides an excellent illustration of our error bounds. For this trajectory, the constant $R$ was calculated at $R=4.218$. Graph (a) shows the true trajectory (in blue) plotted against the forecasted trajectory (in orange). Graph (b) plots the forecast errors (in blue) against the fitted bound (in orange), and Graph (c) is the same, plotted in the log scale. Graph (c) illustrates the similarity between the gradient of the log error bound, given by the top Lyapunov exponent associated to the ESN, and the gradient of the log error growth, demonstrating that, when the ESN has been well-trained, the error bound accurately predicts the rate at which the forecast errors grow.
@@
Figure 5: Prediction error for the Rössler system with full access to system states, averaged over 100 trajectories. The error bound plotted has been averaged over the 100 trajectories. The gray band is an emperically calculated $95\%$ confidence interval for the errors committed by the $100$ trajectories. These are all plotted in the log scale.
To learn the Rössler system, we chose an ESN with $1000$ neurons and connectivity matrix with spectral radius $0.2$. The input mask was scaled by $\gamma=4.65$. Inputs were taken as the states of the entire system, sampled at a time step of $\Delta t=0.005$ over $23000$ time steps. A washout of $2000$ time steps was used. Thereafter, using a Tikhonov regularization, with regularization constant $\alpha=10^{-14.5}$, between the state space trajectory and desired output trajectory on the remaining $21000$ time steps, the readout $h$ was trained and predictions made for an additional $30000$ time steps. Plotted in Figure 5 are the errors in prediction over $6$ Lyapunov times for an RC trained over $100$ trajectories with initial conditions chosen from a Gaussian distribution around $(2,1,5)$ with standard deviation $1\times 10^{-2}$, averaged at each point across the $100$ trajectories and plotted in the log scale. The Lyapunov exponent of the RC was calculated at $\lambda_{1}=0.06933$ using Bennetin’s algorithm over the length of one trajectory. The value of $\delta$ was fixed at $10^{-4}$ and $\varepsilon=9.263\times 10^{-6}$ was calculated by taking the maximum one-step ahead prediction error of the trained ESN over the training portions of the $100$ trajectories. The constant $R$ from (3.12) was calculated for each trajectory using (4.5) and then averaged. Figure 5 shows the error bound using this average value, plotted in the log scale. The log errors grow approximately at a linear rate until saturation of the attractor around $5$ Lyapunov times. The gradient of the linear growth is closely matched by the predicted error bound, whose gradient is given by the top Lyapunov exponent of the RC.
We summarize a few values from the numerical experiments in Table 4.1.
Lorenz full access Lorenz $x$ access Rössler full access
$\lambda_{1}$ $0.8733$ $1.015$ $0.06933$
$\varepsilon$ $8.988\times 10^{-6}$ $5.733\times 10^{-7}$ $9.263\times 10^{-6}$
$L_{z}$ $82.03$ $44.13$ $88.02$
$L_{h}$ $0.2216$ $0.1507$ $0.5116$
$R$ max $2.597$ $883.7$ $15.46$
$R$ mean $0.4089$ $62.22$ $2.497$
$R$ sd $0.01407$ $9.700$ $0.03031$
Table 4.1: Values from the numerical experiments.
@@ Discussion.
The numerical experiments illustrate how the error committed using an ESN to predict three different systems grows at an exponential rate, governed by the top Lyapunov exponent associated with the learned ESN. There is an interplay between the Lyapunov exponents of the dynamical system being learnt and those of the trained ESN. This is seen in its effect on the time until saturation. In the case of the Lorenz attractor, which is known to have a top Lyapunov exponent of $0.9057$ [Visw 98], the top Lyapunov exponent of the ESN trained on the full system states is closer to that of the true system than the top Lyapunov exponent of the ESN trained on a single coordinate. This is to be expected, given that the former ESN is given access to more information during training. Nevertheless, the ESN trained on a single coordinate does remarkably well, on average reaching saturation error around $7$ Lyapunov times, while the ESN trained on full system states saturates on average at $8$ Lyapunov times.
Notably, the ESN trained on the full system states has top Lyapunov exponent lower than that of the true Lorenz system. As discussed in Remark 3.17, decreasing the top Lyapunov exponent of the trained system arbitrarily is expected to lead to an increase in other constants in the bound. The Rössler system has a much lower top Lyapunov exponent than the Lorenz system, and hence in learning it is possible to make accurate predictions significantly further into the future. However, when we standardize the time scale in terms of Lyapunov times, the ESNs trained on the Lorenz system seem to outperform that trained on the Rössler system, which reaches saturation between $5$ and $6$ Lyapunov times on average. The reasons for this may be a combination of a diverse range of factors from the geometry of the different attractors to the care taken in tuning hyperparameters; at the very least this serves to show that the maximum Lyapunov exponent of a system is not the only factor at play. The performance of a single algorithm may vary across different systems.
An extensive discussion on the performance of different algorithms in forecasting chaotic systems is beyond the scope of this paper. Notable progress has been made in [Gilp 23], where the authors propose a database of $135$ chaotic systems and benchmark standard time-series forecasting methods against larger domain-agnostic algorithms on this database. The same author makes further progress in [Zhan 24], where algorithms specialized for the chaotic system forecasting task are compared to more general purpose foundation models performing zero-shot learning given limited training data. Another important study is [Scho 24] where the authors propose a new metric, the ‘cumulative maximum error’, for evaluating the learning performance of different algorithms. They also consider the ‘valid time’ up to which an algorithm makes forecasts whose error remains below a fixed threshold. This is analogous to the time up to saturation in our graphs, putting a well-defined metric on the time up to which an algorithm yields meaningful forecasts.
Notably, ESNs perform well amongst the algorithms they consider. Since ESNs are able to give predictions that remain valid over a long time scale, this makes them well-suited for our purpose of illustrating the error bounds derived in this paper. While the error bound remains valid even for algorithms that do not exhibit good learning, in such cases the linear growth rate in the log errors is not as visually apparent as saturation is quickly reached. This was seen, for example, in trajectories originating from initial conditions for which the ESNs we trained could not make accurate long term predictions. Nevertheless, averaging over a number of randomly generated initial conditions minimizes the effect of initial conditions for which the predictions were not as good, which tend to be a smaller proportion, and shows that in general the linear growth rate of the log errors is indeed matched by the top Lyapunov exponent of the trained system.
It was, however, necessary to draw the initial conditions from a distribution concentrated on a small portion of the attractor. The reason for this was that when training the ESN on initial conditions distributed widely across the attractor, learning was not as good and predictions quickly reached saturation. The error bounds given in this paper find their usefulness precisely when a state-space system has been well-trained and is able to make accurate long-term predictions, since for these systems we may begin to ask further questions, such as at what rate the accuracy of the predictions deteriorates, and up to what time we can trust its predictions. In the case of ill-trained systems, any predictions will soon lose their usefulness. Thus it was necessary to train ESNs that exhibited accurate predictions in the long range.
@@ 5 Conclusions
As noted in the numerical section, the bounds given have a number of limitations. In particular, the theory does not provide an algorithm for calculating a constant $R$ needed for fitting the bound to a given initial condition, nor does it provide a means for calculating the time horizon up to which the bound remains valid. Instead we propose a numerical means of estimating $R$, and find that, empirically, the bound remains valid over the entire trajectory, providing meaningful information until saturation. In the case of well-trained state space systems used in prediction, the bound demonstrates a striking sharpness, accurately predicting the growth rate of the true forecasting error.